OpenStudy (anonymous):

Katniss fires an arrow towards another tribute with a peed of 7.0 x 10^3 m/s. The arrow has a mass of 105 grams and the "dead" tribute has (had) a mass of 58 kg. a. Assume that the arrow became lodged in the tribute; determine the velocity of the arrow tribute system. b. Calculate the kinetic energy of arrow-tribute system

OpenStudy (anonymous):

conservation of momentum: total momentum before = total momentum after \[m _{1}v _{1} + m _{2}v _{2} = m _{1}\prime v _{1}\prime + m _{2}\prime v _{2} \prime \]

OpenStudy (anonymous):

m1=105g, m2=58kg v1=7.0x10^3m/s or is it 0

OpenStudy (anonymous):

For this question, momentum before is just 0.105 x 7000 And after, the total momentum is (0.105 + 58) x velocity of both together

OpenStudy (anonymous):

sorry, you are right, and v1 is 7000, v2 = 0 (the tribute is not moving)

OpenStudy (anonymous):

just follow my method now, do you understand it ok?

OpenStudy (anonymous):

and v2' is still 0 rite?

OpenStudy (anonymous):

i got v1'= 7000m/s o.O

OpenStudy (anonymous):

thats my answer because we want to know v1' for part a.) rite? o:

OpenStudy (anonymous):

this question is a special case of that formula, as the two masses join, and have one velocity. Basically\[v _{1}\prime = v _{2}\prime \]so the right hand side can be simplified to:\[m _{1}\prime v _{1}\prime + m _{2}\prime v _{1}\prime = (m _{1}\prime + m _{2}\prime )v _{1}\prime \]

OpenStudy (anonymous):

oh because the arrow stops right after it hits the tribute rite? o:

OpenStudy (anonymous):

so were looking for vf? o.O

OpenStudy (anonymous):

well, the arrow doesn't stop, it 'takes' the tribute with it. You could just say v1' = v2' = vf. my answer is 12.6 m/s for a)

OpenStudy (anonymous):

so i got 4,509.2

OpenStudy (anonymous):

:/ i think you may not be converting units, everything must be in kg, to get m/s the equation is: 7000 x 0.105 = (58 + 0.105)vf vf = (7000x0.105)/(58+0.105)

OpenStudy (anonymous):

ready for part b), or are you ok with that bit?

OpenStudy (anonymous):

its not .105 its 105 xD anyways thank you~ i got the same answer :DDD

OpenStudy (anonymous):

no i dont get part b.)

OpenStudy (anonymous):

KE=1/2mv^2?

OpenStudy (anonymous):

It is 0.105, 105g = 0.105kg You must always convert units!

OpenStudy (anonymous):

perfect! But what will you put in for m, v?

OpenStudy (anonymous):

oh ok :)

OpenStudy (anonymous):

v=7000 and idk for m? xD

OpenStudy (anonymous):

remember questions are often scaffolded. This means that that you use something you just worked out in the next part. It wants the KE for the system. (the system we just worked out the velocity of!) so v = 12.6 m/s m = (m1 + m2)